# AI-Generated Text is Non-Stationary: Detection via Temporal Tomography

### v1

@@ Abstract
The field of AI-generated text detection has evolved from supervised classification to zero-shot statistical analysis. However, current approaches share a fundamental limitation: they aggregate token-level measurements into scalar scores, discarding positional information about where anomalies occur. Our empirical analysis reveals that AI-generated text exhibits significant non-stationarity—statistical properties vary by 73.8% more between text segments compared to human writing. This discovery explains why existing detectors fail against localized adversarial perturbations that exploit this overlooked characteristic. We introduce Temporal Discrepancy Tomography (TDT), a novel detection paradigm that preserves positional information by reformulating detection as a signal processing task. TDT treats token-level discrepancies as a time-series signal and applies Continuous Wavelet Transform to generate a two-dimensional time-scale representation, capturing both the location and linguistic scale of statistical anomalies. On the RAID benchmark, TDT achieves 0.855 AUROC (7.1% improvement over the best baseline). More importantly, TDT demonstrates robust performance on adversarial tasks, with 14.1% AUROC improvement on HART Level 2 paraphrasing attacks. Despite its sophisticated analysis, TDT maintains practical efficiency with only 13% computational overhead. Our work establishes non-stationarity as a fundamental characteristic of AI-generated text and demonstrates that preserving temporal dynamics is essential for robust detection.
= Date: August 4, 2025
[Uncaptioned image] Code Repository: https://github.com/ResearAI/TDT-Text-Detect
= Contact: zhangyue@westlake.edu.cn
@@ 1 Introduction
The widespread deployment of large language models has fundamentally altered the landscape of content creation, from academic writing to journalism and social media. This transformation brings unprecedented challenges for maintaining information integrity, as distinguishing between human and machine-generated text becomes increasingly difficult yet critically important (Jawahar et al., 2020). The sophistication of modern language models enables not only wholesale generation of convincing text but also subtle modifications that preserve human-like qualities while introducing machine artifacts (Su et al., 2025, Zhang et al., 2024).
@@
Extensive evaluation validates our approach. TDT achieves 0.855 AUROC on the RAID benchmark (7.1% improvement) and excels on adversarial tasks with 14.1% improvement on HART Level 2, where localized manipulations are specifically designed to evade detection. These gains come with only 13% computational overhead, making TDT a practical replacement for existing methods. Our contributions are threefold. We provide the first empirical evidence that non-stationarity is a fundamental characteristic of AI-generated text, a property unaddressed by current methods. We then demonstrate that by preserving this positional information through signal processing techniques, detection robustness is significantly improved, particularly against adversarial attacks. Finally, based on these insights, we establish a new detection paradigm that analyzes temporal dynamics to achieve state-of-the-art performance efficiently, charting a more robust path forward for the field.
@@ 2 Related Work
Beyond purely statistical zero-shot methods, the detection landscape includes other important paradigms. Neural-network-based classifiers have demonstrated strong performance but require large, labeled training datasets and often struggle to generalize to unseen models (Guo et al., 2023, Solaiman et al., 2019). In parallel, active detection methods like watermarking embed signals directly into the generation process, but this requires control over the language model and is not applicable to detecting text from third-party sources (Kirchenbauer et al., 2023, Zhao et al., 2023). Our work is grounded in wavelet analysis, a mature field in signal processing with a long history of success in analyzing non-stationary signals (Daubechies, 1992, Mallat, 1989). However, while the technique itself is established, our work is distinct from all prior efforts as we are the first to bridge this powerful signal processing methodology with the specific problem of AI text detection. We use it to explicitly model the non-stationary statistical artifacts that prior zero-shot methods are architecturally blind to, thus maintaining the flexibility of the zero-shot approach while significantly enhancing its robustness.
@@ 3 Method
The central premise of our work is that the location of statistical anomalies within a text is as important as their magnitude. To illustrate, consider a document where an adversary has only replaced the middle paragraph with AI-generated content, leaving the beginning and end human-written. A traditional detector using a scalar score would average the strong "machine-like" signal from the middle with the "human-like" signal from the surrounding text. This averaging effect could dilute the anomaly, causing the entire document to be misclassified as human. Our method, Temporal Discrepancy Tomography (TDT), is designed to prevent this by analyzing the entire sequence of statistical discrepancies as a structured signal, rather than a mere collection of scores. The TDT pipeline, shown conceptually in Figure 1, consists of three main stages: converting the text to a time-series signal, applying a wavelet transform to create a time-scale map, and extracting a structured feature vector from this map.
@@ 3.1 Step 1: From Text to a Time-Series Signal
The TDT pipeline begins with a sequence of token-level discrepancy scores, $Z(x)=[z_{1},z_{2},...,z_{n}]$. Each score, $z_{i}$, quantifies the statistical "surprise" of the $i$-th token. For this, we adopt the robust t-distribution normalization from the T-Detect framework (West et al., 2025). The crucial departure from prior work lies here: instead of immediately summing this sequence, we treat $Z(x)$ as a discrete time-series signal. This shift in perspective is the foundation of our method. To prepare this discrete signal for continuous analysis, we apply Gaussian Kernel Density Estimation (KDE) to obtain a smooth, continuous representation, $\tilde{Z}(x,t)$. This is a standard signal processing step that allows the application of techniques like the Continuous Wavelet Transform while preserving the underlying structure of the token-level data (Elouaham et al., 2024, Noskova and Tumakov, 2024). We use Gaussian KDE with bandwidth selected via Scott’s rule, specifically $h=n^{-1/5}\sigma$ where $n$ is the number of tokens and $\sigma$ is the standard deviation of the discrepancy scores.
@@ 3.2 Step 2: Wavelet Transform for Time-Scale Analysis
[…]
Here, the translation parameter $b$ slides the wavelet $\psi$ across the signal, telling us where in the text we are looking. Where $\psi^{*}$ denotes the complex conjugate of the mother wavelet $\psi$. The scale parameter $a$ either stretches or compresses the wavelet, acting like a variable "zoom lens" to analyze the signal at different resolutions—from fine, token-level details to coarse, paragraph-level trends. Based on extensive ablation studies, we selected the Complex Morlet wavelet ($\psi(t)=\pi^{-1/4}e^{i\omega_{0}t}e^{-t^{2}/2}$ with $\omega_{0}=6$), prized for its excellent trade-off between time and frequency localization (Mohamed et al., 2023). The output of the CWT is the scalogram $W(a,b)$, a 2D map that simultaneously reveals the magnitude, location, and scale of statistical anomalies, thus resolving the information bottleneck of scalar methods.
3.3 Step 3: Hierarchical Feature Extraction
This entire feature extraction process adds only a modest 13% latency overhead compared to its scalar counterpart, making TDT a practical, powerful, and more informative "drop-in replacement" for the summarization step in existing detection pipelines.
4 Experimental Setup
To ensure a fair and rigorous comparison, all discrepancy-based methods, including our proposed TDT, utilize the same core model architecture. We use the high-performing Falcon-7B as the reference model and Falcon-7B-Instruct as the scoring model, following established practices that have demonstrated their effectiveness in generating the statistical artifacts central to this detection paradigm (West et al., 2025). The Binoculars baseline is evaluated using its standard, publicly available configuration (with Falcon-7B and Falcon-7B-Instruct). All input texts are truncated to a maximum of 512 tokens. Our evaluation spans a suite of diverse benchmarks: the adversarial RAID benchmark (Dugan et al., 2024), which tests robustness against various manipulation techniques; the multi-level HART benchmark (Reid et al., 2024), which assesses performance on simple detection, adversarial paraphrasing, and humanization; and for generalization, we use text from the architecturally distinct QWEN-3-0.6B model and non-English news domains (Spanish and Arabic).
@@ 4 Experimental Setup
Level 1 (Simple Detection) Level 2 (Adversarial Paraphrasing) Level 3 (Humanization) Overall Results
Method Essay News Writing Arxiv Essay News Writing Arxiv Essay News Writing Arxiv L1 L2 L3
FastDetectGPT 0.877 0.714 0.740 0.769 0.734 0.689 0.692 0.718 0.883 0.851 0.840 0.877 0.778 0.711 0.862
Binoculars 0.879 0.720 0.740 0.769 0.735 0.699 0.693 0.715 0.897 0.866 0.847 0.882 0.780 0.711 0.870
T-Detect 0.880 0.714 0.740 0.771 0.734 0.698 0.693 0.718 0.891 0.863 0.844 0.879 0.780 0.712 0.867
TDT (Ours) 0.882 0.778 0.815 0.828 0.746 0.815 0.842 0.858 0.890 0.869 0.900 0.919 0.825 0.812 0.891
$\Delta$ vs Best +0.2% +8.1% +10.1% +7.4% +1.5% +16.7% +21.5% +19.5% -0.8% +0.7% +6.3% +4.5% +5.8% +14.1% +2.4%
@@ 5.1 Main Performance Evaluation
This trend of robust performance is further confirmed on the HART benchmark. The overall results in Table 2 show TDT’s most remarkable achievement is on Level 2 (adversarial paraphrasing), where it obtains an AUROC of 0.812—a dramatic 14.1% improvement over all baselines. The domain-specific results reveal that this gain is driven by exceptional performance on domains like Writing (+21.5%) and Arxiv (+19.5%). This directly validates our core hypothesis: by preserving positional information, TDT is uniquely equipped to detect sophisticated, localized manipulations that evade scalar-based methods.
Level 1 (Simple Detection) Level 2 (Adversarial Paraphrasing) Level 3 (Humanization) Overall Results
Method Essay News Writing Arxiv Essay News Writing Arxiv Essay News Writing Arxiv L1 L2 L3
FastDetectGPT 0.589 0.579 0.601 0.647 0.443 0.563 0.674 0.461 0.633 0.632 0.601 0.613 0.448 0.505 0.645
Binoculars 0.589 0.580 0.601 0.647 0.436 0.556 0.674 0.454 0.649 0.639 0.601 0.635 0.446 0.504 0.650
T-Detect 0.590 0.582 0.601 0.642 0.440 0.557 0.674 0.463 0.632 0.637 0.601 0.618 0.448 0.510 0.642
TDT (Ours) 0.601 0.642 0.601 0.712 0.674 0.699 0.674 0.652 0.537 0.673 0.601 0.623 0.629 0.668 0.607
$\Delta$ vs Best +1.9% +10.3% 0.0% +10.0% +53.0% +25.5% 0.0% +40.8% -17.3% +5.4% 0.0% -1.9% +40.4% +31.0% -6.6%
Table 3: HART Benchmark performance (AUROC) on QWEN-3-0.6B results, presented in a stacked format. The ’Overall Results’ section shows the summary scores for each level.
@@ 5.2 Analysis through Research Questions
5.2.1 RQ1: How can the information loss from scalar summarization be overcome, and does AI text exhibit the non-stationary properties that make this necessary?
@@
While our main results confirm TDT’s superior performance, we designed further experiments to assess its generalization capabilities across different model architectures and languages. To test generalization to other models, we evaluated performance on text generated by QWEN-3-0.6B. The experimental phenomena, detailed in Tables 3, 4, and 5, show that TDT’s advantages are not confined to a single setup. On the English RAID domains, TDT achieves an overall AUROC of 0.724, a 6.3% improvement over the best baseline (Table 4). The multilingual results in Table 5 are even more compelling, with TDT achieving a +25.5% AUROC gain on Spanish text and a +40.8% gain on Arabic text for HART Level 2.

### v2

@@ Abstract
The field of AI-generated text detection has evolved from supervised classification to zero-shot statistical analysis. However, current approaches share a fundamental limitation: they aggregate token-level measurements into scalar scores, discarding positional information about where anomalies occur. Our empirical analysis reveals that AI-generated text exhibits significant non-stationarity—statistical properties vary by 73.8% more between text segments compared to human writing. This discovery explains why existing detectors fail against localized adversarial perturbations that exploit this overlooked characteristic. We introduce Temporal Discrepancy Tomography (TDT), a novel detection paradigm that preserves positional information by reformulating detection as a signal processing task. TDT treats token-level discrepancies as a time-series signal and applies Continuous Wavelet Transform to generate a two-dimensional time-scale representation, capturing both the location and linguistic scale of statistical anomalies. On the RAID benchmark, TDT achieves 0.855 AUROC (7.1% improvement over the best baseline). More importantly, TDT demonstrates robust performance on adversarial tasks, with 14.1% AUROC improvement on HART Level 2 paraphrasing attacks. Despite its sophisticated analysis, TDT maintains practical efficiency with only 13% computational overhead. Our work establishes non-stationarity as a fundamental characteristic of AI-generated text and demonstrates that preserving temporal dynamics is essential for robust detection. Ours code are released at https://github.com/ResearAI/TDT-Text-Detect.
WARMING: We hereby declare that ours DeepScientist system performed approximately 95% of the work presented in this paper. This includes the initial ideation, the design and execution of comparative experiments, the analysis of results, the literature review, the composition of the manuscript, the creation of the main figures, and the organization of the accompanying code repository. The role of the human authors was to supervise the AI’s operations. While we have diligently worked to minimize AI hallucinations and ensure the validity of the experimental results, we cannot fully guarantee against potential unintended outputs, system failures, or misleading conclusions. We therefore advise readers to approach this work with caution and to critically evaluate its findings before application.
@@ 1 Introduction
The widespread deployment of large language models has fundamentally altered the landscape of content creation, from academic writing to journalism and social media. This transformation brings unprecedented challenges for maintaining information integrity, as distinguishing between human and machine-generated text becomes increasingly difficult yet critically important (Jawahar et al., 2020). The sophistication of modern language models enables not only wholesale generation of convincing text but also subtle modifications that preserve human-like qualities while introducing machine artifacts (Su et al., 2025; Zhang et al., 2024).
@@
Extensive evaluation validates our approach. TDT achieves 0.855 AUROC on the RAID benchmark (7.1% improvement) and excels on adversarial tasks with 14.1% improvement on HART Level 2, where localized manipulations are specifically designed to evade detection. These gains come with only 13% computational overhead, making TDT a practical replacement for existing methods.
Our contributions are threefold:
  • We provide empirical evidence that non-stationarity is a fundamental characteristic of AI-generated text, not captured by current detection methods.
  • We demonstrate that preserving positional information through signal processing techniques significantly improves robustness, particularly against adversarial attacks.
  • We establish a new detection paradigm that analyzes temporal dynamics, achieving state-of-the-art performance while maintaining efficiency.
@@ 2 Related Work
Beyond purely statistical zero-shot methods, the detection landscape includes other important paradigms. Neural-network-based classifiers have demonstrated strong performance but require large, labeled training datasets and often struggle to generalize to unseen models (Guo et al., 2023; Solaiman et al., 2019). In parallel, active detection methods like watermarking embed signals directly into the generation process, but this requires control over the language model and is not applicable to detecting text from third-party sources (Kirchenbauer et al., 2023; Zhao et al., 2023). Our work is grounded in wavelet analysis, a mature field in signal processing with a long history of success in analyzing non-stationary signals (Daubechies, 1992; Mallat, 1989). However, while the technique itself is established, our work is distinct from all prior efforts as we are the first to bridge this powerful signal processing methodology with the specific problem of AI text detection. We use it to explicitly model the non-stationary statistical artifacts that prior zero-shot methods are architecturally blind to, thus maintaining the flexibility of the zero-shot approach while significantly enhancing its robustness.
@@ 3 Method
The central premise of our work is that the location of statistical anomalies within a text is as important as their magnitude. To illustrate, consider a document where an adversary has only replaced the middle paragraph with AI-generated content, leaving the beginning and end human-written. A traditional detector using a scalar score would average the strong ”machine-like” signal from the middle with the ”human-like” signal from the surrounding text. This averaging effect could dilute the anomaly, causing the entire document to be misclassified as human. Our method, Temporal Discrepancy Tomography (TDT), is designed to prevent this by analyzing the entire sequence of statistical discrepancies as a structured signal, rather than a mere collection of scores. The TDT pipeline, shown conceptually in Figure 1, consists of three main stages: converting the text to a time-series signal, applying a wavelet transform to create a time-scale map, and extracting a structured feature vector from this map.
@@ 3.1 Step 1: From Text to a Time-Series Signal
The TDT pipeline begins with a sequence of token-level discrepancy scores, $Z(x)=[z_{1},z_{2},...,z_{n}]$. Each score, $z_{i}$, quantifies the statistical ”surprise” of the $i$-th token. For this, we adopt the robust t-distribution normalization from the T-Detect framework (West et al., 2025). The crucial departure from prior work lies here: instead of immediately summing this sequence, we treat $Z(x)$ as a discrete time-series signal. This shift in perspective is the foundation of our method. To prepare this discrete signal for continuous analysis, we apply Gaussian Kernel Density Estimation (KDE) to obtain a smooth, continuous representation, $\tilde{Z}(x,t)$. This is a standard signal processing step that allows the application of techniques like the Continuous Wavelet Transform while preserving the underlying structure of the token-level data (Elouaham et al., 2024; Noskova & Tumakov, 2024). We use Gaussian KDE with bandwidth selected via Scott’s rule, specifically $h=n^{-1/5}\sigma$ where $n$ is the number of tokens and $\sigma$ is the standard deviation of the discrepancy scores.
@@ 3.2 Step 2: Wavelet Transform for Time-Scale Analysis
$W(a,b)=\frac{1}{\sqrt{a}}\int_{-\infty}^{\infty}\tilde{Z}(x,t)\psi^{*}\left(\frac{t-b}{a}\right)dt$ (1)
Here, the translation parameter $b$ slides the wavelet $\psi$ across the signal, telling us where in the text we are looking. Where $\psi^{*}$ denotes the complex conjugate of the mother wavelet $\psi$. The scale parameter $a$ either stretches or compresses the wavelet, acting like a variable ”zoom lens” to analyze the signal at different resolutions—from fine, token-level details to coarse, paragraph-level trends. Based on extensive ablation studies, we selected the Complex Morlet wavelet ($\psi(t)=\pi^{-1/4}e^{i\omega_{0}t}e^{-t^{2}/2}$ with $\omega_{0}=6$), prized for its excellent trade-off between time and frequency localization (Mohamed et al., 2023). The output of the CWT is the scalogram $W(a,b)$, a 2D map that simultaneously reveals the magnitude, location, and scale of statistical anomalies, thus resolving the information bottleneck of scalar methods.
@@
3.3 Step 3: Hierarchical Feature Extraction
This entire feature extraction process adds only a modest 13% latency overhead compared to its scalar counterpart, making TDT a practical, powerful, and more informative ”drop-in replacement” for the summarization step in existing detection pipelines.
4 Experimental Setup
Overall Results (AUROC)
Method L1 L2 L3
FastDetectGPT 0.778 0.711 0.862
Binoculars 0.780 0.711 0.870
T-Detect 0.780 0.712 0.867
TDT (Ours) 0.825 0.812 0.891
$\Delta$ vs Best +5.8% +14.1% +2.4%
Table 2: Overall performance (AUROC) on the HART Benchmark.
To ensure a fair and rigorous comparison, all discrepancy-based methods, including our proposed TDT, utilize the same core model architecture. We use the high-performing Falcon-7B as the reference model and Falcon-7B-Instruct as the scoring model, following established practices that have demonstrated their effectiveness in generating the statistical artifacts central to this detection paradigm (West et al., 2025). The Binoculars baseline is evaluated using its standard, publicly available configuration (with Falcon-7B and Falcon-7B-Instruct). All input texts are truncated to a maximum of 512 tokens. Our evaluation spans a suite of diverse benchmarks: the adversarial RAID benchmark (Dugan et al., 2024), which tests robustness against various manipulation techniques; the multi-level HART benchmark (Bao et al., 2025), which assesses performance on simple detection, adversarial paraphrasing, and humanization; and for generalization, we use text from the architecturally distinct QWEN-3-0.6B model and non-English news domains (Spanish and Arabic).
@@ 5 Experiments and Results
Level 1 (Simple Detection)
Method Essay News Writing Arxiv
F-D-GPT 0.877 0.714 0.740 0.769
Binoculars 0.879 0.720 0.740 0.769
T-Detect 0.880 0.714 0.740 0.771
TDT (Ours) 0.882 0.778 0.815 0.828
$\Delta$ vs Best +0.2% +8.1% +10.1% +7.4%
Level 2 (Adversarial Paraphrasing)
Method Essay News Writing Arxiv
F-D-GPT 0.734 0.689 0.692 0.718
Binoculars 0.735 0.699 0.693 0.715
T-Detect 0.734 0.698 0.693 0.718
TDT (Ours) 0.746 0.815 0.842 0.858
$\Delta$ vs Best +1.5% +16.7% +21.5% +19.5%
Level 3 (Humanization)
Method Essay News Writing Arxiv
F-D-GPT 0.883 0.851 0.840 0.877
Binoculars 0.897 0.866 0.847 0.882
T-Detect 0.891 0.863 0.844 0.879
TDT (Ours) 0.890 0.869 0.900 0.919
$\Delta$ vs Best -0.8% +0.3% +6.3% +4.2%
@@
This trend of robust performance is further confirmed on the HART benchmark (Bao et al., 2025). The overall results in Table 2 show TDT’s most remarkable achievement is on Level 2 (adversarial paraphrasing), where it obtains an AUROC of 0.812—a dramatic 14.1% improvement over all baselines. The domain-specific results in Table 3 reveal that this gain is driven by exceptional performance on domains like Writing (+21.5%) and Arxiv (+19.5%). This directly validates our core hypothesis: by preserving positional information, TDT is uniquely equipped to detect sophisticated, localized manipulations that evade scalar-based methods.
@@ 5.1 Analysis through Research Questions
5.1.1 RQ1: How can the information loss from scalar summarization be overcome?
@@
Level 1 (Simple Detection)
Method Essay News Writing Arxiv
F-DetectGPT 0.589 0.579 0.601 0.647
Binoculars 0.589 0.580 0.601 0.647
T-Detect 0.590 0.582 0.601 0.642
TDT (Ours) 0.601 0.642 0.601 0.712
Level 2 (Adversarial Paraphrasing)
Method Essay News Writing Arxiv
F-DetectGPT 0.443 0.563 0.674 0.461
Binoculars 0.436 0.556 0.674 0.454
T-Detect 0.440 0.557 0.674 0.463
TDT (Ours) 0.674 0.699 0.674 0.652
Level 3 (Humanization)
Method Essay News Writing Arxiv
F-DetectGPT 0.633 0.632 0.601 0.613
Binoculars 0.649 0.639 0.601 0.635
T-Detect 0.632 0.637 0.601 0.618
TDT (Ours) 0.537 0.673 0.601 0.623
Table 6: HART Benchmark performance (AUROC) on QWEN-3-0.6B results.
@@
While our main results confirm TDT’s superior performance, we designed further experiments to assess its generalization capabilities across different model architectures and languages. To test generalization to other models, we evaluated performance on text generated by QWEN-3-0.6B. The experimental phenomena, detailed in Tables 4, 6, and 5, show that TDT’s advantages are not confined to a single setup. On the English RAID domains, TDT achieves an overall AUROC of 0.724, a 6.3% improvement over the best baseline (Table 4). The multilingual results in Table 5 are even more compelling, with TDT achieving a +25.5% AUROC gain on Spanish text and a +40.8% gain on Arabic text for HART Level 2.